# Align & Invert: Solving Inverse Problems with Diffusion and Flow-based Models via Representational Alignment

### v1

@@ Abstract
Enforcing alignment between the internal representations of diffusion or flow-based generative models and those of pretrained self-supervised encoders has recently been shown to provide a powerful inductive bias, improving both convergence and sample quality. In this work, we extend this idea to inverse problems, where pretrained generative models are employed as priors. We propose applying representation alignment (Repa) between diffusion or flow-based models and a pretrained self-supervised visual encoder, such as DINOv2, to guide the reconstruction process at inference time. Although ground-truth signals are unavailable in inverse problems, we show that aligning model representations with approximate target features can substantially enhance reconstruction fidelity and perceptual realism. We provide theoretical results showing (a) the relation between the Repa regularization and a divergence measure in the DINOv2 embedding space, and (b) how Repa updates steer the model’s internal representations toward those of the clean image. These results offer insights into the role of Repa in improving perceptual fidelity. Finally, we demonstrate the generality of our approach by integrating it into multiple state-of-the-art inverse problem solvers. Extensive experiments on super-resolution, box inpainting, Gaussian deblurring, and motion deblurring confirm that our method consistently improves reconstruction quality across tasks, while also providing substantial efficiency gains reducing the number of required discretization steps without compromising the performance of the underlying solver. Our code is available at https://github.com/Sfountouris/Align_And_Invert.
@@ 1 Introduction
Pretrained diffusion and flow-based models, (sohl2015deep; ho2020denoising; esser2024scaling; lipman2023flow), have recently been at the heart of methods focusing on addressing various types of inverse problems. The crux of these approaches is to perform diffusion sampling, incorporating measurement information during the reverse process, with the goal to reconstruct the clean image at the final step, (patel2024steering; chung2023diffusion). Undoubtedly, diffusion and flow-based models have pushed the boundaries in addressing a wealth of inverse problems, (daras2024survey). However, they still struggle to produce highly detailed images particularly in cases of severe degradation and in challenging scenes, such as natural images with rich textures and complex backgrounds. These issues are further exacerbated when using latent diffusion models (rombach2022high), where the inclusion of the encoder-decoder architecture introduces additional challenges during the reverse process due to the nonlinearities introduced. To mitigate these limitations, further inductive biases are needed during inference, (rout2023solving).
@@
Measurement Without REPA With REPA Ground Truth Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionDINOv2 EncoderRefer to caption Approximate reconstructed imageRepresentation AlignmentObjectiveSiT Transformer BlockSiT Transformer BlockSiT Transformer BlockSiT Transformer BlockSiT Transformer BlockSiT Transformer BlockMeasurement Matching ObjectiveMLP+
@@
Figure 1: Overview of our proposed framework. Left: Box inpainting (top row) and Gaussian deblurring (bottom row) results, where adding Repa (third column) improves the perceptual quality of reconstructed images compared to the baseline without Repa (second column). Right: Alignment mechanism between diffusion-model features and pretrained DINOv2 embeddings, combined with measurement matching. An approximate reconstructed image is provided as input to the DINOv2 encoder for computing the alignment signal.
@@
Recent studies have shown that diffusion models learn semantic features in their hidden states, (li2023your). The more expressive these features are the better the diffusion model performs of on the generative task, (xiang2023). Building on this insight, the seminal work of yu2024repa introduced a regularizer that aligns the internal representations of the diffusion model with those of a pretrained visual encoder, (oquab2023dinov2). This framework, termed representation alignment, was shown to act as a strong semantic constraint leading to higher-fidelity generations and significantly faster convergence.
@@
The success of Repa has attracted considerable interest in this direction, (wang2025repa; tian2025u; yao2025reconstruction; leng2025repa; wang2025learning). In particular, (wang2025repa) observed that while Repa enhances early training dynamics, its benefits diminish over time and can even degrade performance in later iterations. To address this issue, they introduced an early-stopping strategy which mitigated the late-stage degradation and further improved overall convergence speed. Other works have explored applying alignment losses directly to VAE latent spaces (yao2025reconstruction; xu2025exploring), finding that diffusion models trained on these regularized spaces converge faster. Despite these advances, most efforts focus on improving training efficiency, whereas relatively little attention has been paid to leveraging representation alignment during inference, particularly for solving inverse problems. This motivates our central question:
@@
Can we apply representation alignment ideas to benefit existing algorithms for solving inverse problems using pretrained diffusion and flow-based models?
@@
We introduce a general framework for solving inverse problems with diffusion and flow-based models by enforcing alignment between internal diffusion representations and DINOv2 features. Despite the absence of ground truth, alignment with proxy reconstructions remains effective thanks to the robustness of DINOv2 features (see Figure 1). We provide theoretical results that demonstrate perceptual benefits of Repa regularizer showing a) its relation with a divergence measure defined on the DINOv2 feature space and and b) how it enhances convergence of the model’s internal representations toward those of the clean image. We demonstrate the versatility of our approach by integrating representational alignment into different existing state-of-the-art inverse problem solvers. Through extensive experiments on super-resolution, Gaussian deblurring, motion deblurring, and box inpainting, we show consistent improvements over prior methods across tasks and metrics, and further demonstrate that applying our regularizer to existing solvers allows them to match their original performance using significantly fewer discretization steps.
@@ 2.1 Diffusion & Flow-based Models
Generative models based on denoising are a class of deep generative models that aim to learn a data distribution $p_{0}(x)$ by training on progressively corrupted versions of real samples (sohl2015deep; ho2020denoising; song2021denoising; liu2023flow). These models define a) a forward (noising) process, which gradually interpolates between the data distribution and a reference distribution $p_{1}(x)$ (typically a Gaussian), and b) a reverse (denoising) process, which is trained to invert this corruption and iteratively reconstruct data from noise. Following the formulation of stochastic interpolants (albergo2023stochastic), the time marginals of this process can be expressed as
@@
where $\alpha_{t}$ and $\sigma_{t}$ correspond to decreasing and increasing function of $t$, respectively, with boundary conditions $\alpha_{0}=\sigma_{1}=1$ and $\alpha_{1}=\sigma_{0}=0$. Throughout, we use the notation $\dot{\alpha}_{t}:=\tfrac{\mathrm{d}\alpha_{t}}{\mathrm{d}t}$ and $\dot{\sigma}_{t}:=\tfrac{\mathrm{d}\sigma_{t}}{\mathrm{d}t}$ to denote time derivatives. lipman2023flow showed that there exists a probability flow ordinary differential equation (ODE) whose solution follows the same time marginals as the process in eq. (1):
@@
where $v(x_{t},t)$ denotes the true drift field that transports $p_{1}(x)$ to $p_{0}(x)$. To enable sampling from the data distribution, one has to train a neural network $v_{\theta}(x_{t},t)$ to approximate $v(x_{t},t)$. This is done by regressing the network output toward the optimal vector field using conditional flow matching (lipman2023flow; hyvarinen2005estimation), minimizing
@@
eq. (5) hints that the score parameterization used in diffusion models and the velocity parameterization used in flow-based models can lead to equivalent formulations of the sampling process. In practice, diffusion models correspond to a stochastic discretization of the probability flow ODE, while flow-based models correspond to its deterministic counterpart.
@@ 2.2 Diffusion Models with Representation Guidance
Diffusion models learn rich, discriminative features in their hidden states, which are crucial for their generative performance (xiang2023). Nevertheless, their learned representations still fall behind on downstream tasks when compared with state-of-the-art self-supervised visual encoders, (assran2023self; simeoni2025dinov3). This representation gap has been identified as a key bottleneck for improving the generative performance of diffusion models, as richer and more semantically structured representations have been shown to enhance sample fidelity (yu2024repa). To address this, a growing line of work introduces representation guidance, which conditions diffusion models on external representations from pretrained self-supervised encoders (li2024return). Typically these approaches fall into two categories: (i) conditioning the diffusion model on external representations, either learned, (li2024return), or retrieved from a database of samples, (blattmann2022retrieval; hu2023self; sheynin2022knn); and (ii) explicitly aligning the model’s internal representations with those of large pretrained encoders, thereby encouraging more semantically meaningful feature-spaces, (tian2025u; hu2023self; leng2025repa).
Interestingly, wang2025learning demonstrated that methods from both families can be unified under a common variational framework by deriving an evidence lower bound (ELBO) conditioned on the auxiliary representation. The main difference between these two families lies in how the auxiliary representations are integrated into the training process: in the first family, the representations are generated independently of the diffusion dynamics and serve solely as conditioning signals, whereas in the second family, the training process itself is explicitly regularized to align its intermediate features with the target representations. In particular, yu2024repa introduced Repa, a method that enforces feature-level alignment in a transformer-based diffusion model (ma2024sit), to improve sample quality and training efficiency by encouraging semantically meaningful representations to emerge inside the diffusion model.
@@
$h_{t}=\textsc{DiffEnc}(x,t),$
where DiffEnc returns the patch-level hidden representation after the first $\ell$ blocks of the diffusion model’s transformer encoder. Repa enforces patch-wise similarity between the $D_{1}$-dimensional DINOv2 representation of the target image $x$, i.e., $f^{[n]}_{\text{DINOv2}}(x)\in\mathbb{R}^{D_{1}}$ for patches $n=1,\dots,N$, and the intermediate hidden states $h^{[n]}_{t}$, using the following loss:
@@
2.3 Solving Inverse Problems With Diffusion and Flow-based Models
When solving inverse problems the goal is to recover the clean signal $x_{0}$ from noisy or degraded measurements $y$. In the flow- and diffusion-based framework, this is accomplished by replacing the unconditional score function in eq. (4) with the conditional score $\nabla_{x_{t}}\log p(x_{t}\mid y)$. Applying Bayes’ rule, the reverse-time SDE becomes
$dx_{t}=\Bigl(v(x_{t},t)-g^{2}(t)\bigl[\nabla_{x_{t}}\log p_{t}(x_{t})+\nabla_{x_{t}}\log p(y\mid x_{t})\bigr]\Bigr)dt+g(t)d\bar{w}_{t},$ (7)
The term $\nabla_{x_{t}}\log p_{t}(x_{t})$ is the unconditional score and can be computed using a pretrained diffusion/flow model. Intuitively, this term guides $x_{t}$ toward the learned data distribution. The term $\nabla_{x_{t}}\log p(y\mid x_{t})$ enforces consistency with the measurements. However, computing $p(y\mid x_{t})$ is generally intractable, as it requires marginalizing over all possible clean signals $x_{0}$.
As a result, practical algorithms rely on approximations of $\nabla_{x_{t}}\log p(y\mid x_{t})$. A recent survey by daras2024survey provides a comprehensive taxonomy of these methods. One of the most prominent approaches is Diffusion Posterior Sampling (DPS), (chung2023diffusion), which uses the following approximation:
$p(y\mid x_{t})=\int p(y\mid x_{0})p(x_{0}\mid x_{t})\,dx_{0}\approx p\left(y\mid\hat{x}_{0}=\mathbb{E}[x_{0}\mid x_{t}]\right),$ (8)
where $\mathbb{E}[x_{0}\mid x_{t}]$ denotes the conditional expectation of $x_{0}$ given $x_{t}$. In many practical scenarios, performing the diffusion process directly in pixel space is computationally prohibitive. Latent diffusion models (LDMs) address this limitation by operating in a compressed latent space. Let $\mathcal{E}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ be an encoder mapping an image $x_{0}$ to its latent representation $z_{0}=\mathcal{E}(x_{0})$, and let $\mathcal{D}$ denote the corresponding decoder such that $\hat{x}_{0}=\mathcal{D}(z_{0})\approx x_{0}$. Building on this formulation, recent work has proposed latent-space counterparts of pixel-based inverse problem solvers. For example, Diffusion Posterior Sampling (DPS) has been formulated in latent space by redefining the measurement likelihood as
$p(y\mid z_{t})\approx p\!\left(y\,\middle|\,\hat{x}_{0}=\mathcal{D}\!\big(\mathbb{E}[z_{0}\mid z_{t}]\big)\right).$ (9)
This straightforward formulation commonly referred to as Latent DPS has been observed to introduce artifacts, largely due to the nonlinearity of the decoder. To mitigate this effect, rout2023solving proposed a regularization term encouraging the latent variables to remain close to a fixed point of the encoder–decoder mapping, while song2024solving enforced data consistency during optimization followed by a projection of the latent variables back onto the noisy manifold.
@@ 3 Proposed Approach
Given the success of the Repa framework, we aim to apply it to inverse problems. However, a key challenge is that representation alignment requires access to the ground-truth signal, which is not available in this setting. To overcome this, we must choose an alternative input to the DINOv2 encoder that can stand in for the missing ground truth. In particular, we need an approximation $\hat{x}_{0}$ of the unknown $x_{0}$ that produces a proxy representation $c_{\text{proxy}}\in\mathbb{R}^{N\times D_{1}}$, where $N$ denotes the number of patch tokens and $D_{1}$ is their feature dimension. This representation should satisfy
$c_{\text{proxy}}\equiv f_{\text{DINOv2}}(\hat{x}_{0})\approx f_{\text{DINOv2}}(x_{0}).$ (10)
Then, we sample using a regularized objective that augments the standard posterior with an additional representation-guidance term:
$J_{t}(x_{t}\mid y,c_{\text{proxy}})=\underbrace{\log p(y\mid x_{t})}_{\text{likelihood}}+\underbrace{\log p_{t}(x_{t})}_{\text{prior}}+\lambda\,\underbrace{\log p(c_{\text{proxy}}\mid x_{t})}_{\text{REPA term}},$ (11)
where $\lambda>0$ controls the strength of the additional guidance term. Here, the first term enforces measurement consistency, the second term corresponds to the unconditional prior and the last term acts as an additional semantic regularizer encouraging alignment with $c_{\text{proxy}}$ in the DINOv2 feature space.
Representation Alignment term. Let us express the flow- or diffusion-based model as $u_{\theta}=G_{1}\circ(G_{2}(x_{t},t))$, and denote $h_{t}^{[n]}=G_{2}^{[n]}(x_{t},t)$ the $n$-th patch feature extracted from the diffusion model’s hidden state $h_{t}$ at timestep $t$. Following the training loss of Repa in eq. (6), we approximate the gradient of the representation likelihood term with a sum of cosine similarities between patch embeddings:
$\nabla_{x_{t}}\log p(c_{\text{proxy}}\mid x_{t})\;\approx\;\nabla_{x_{t}}\sum_{n=1}^{N}\cos\!\bigl(c^{[n]}_{\text{proxy}},\,g_{\phi}(G_{2}^{[n]}(x_{t},t))\bigr),$ (12)
where $g_{\phi}$ is a projection head mapping the model’s features into the DINOv2 feature-space (Section 2.2). When incorporated into the sampling objective, the term in eq. (12) encourages the intermediate representations of the model to align with $c_{\text{proxy}}$, i.e., the DINOv2 representation of $\hat{x}_{0}$ .
@@
On the selection of $c_{\text{proxy}}$. To address the fact that representation alignment (yu2024repa) is not directly applicable to inverse problems where the ground-truth image is unavailable at inference time, we construct $c_{\text{proxy}}$ from an approximate reconstruction. When the measurement $y$ lies in the image domain, we simply use $f_{\text{DINOv2}}(y)$. Otherwise, we form a proxy using the best denoised estimate available at timestep $t$, namely $f_{\text{DINOv2}}(\mathbb{E}[x_{0}\mid x_{t}])$. In both cases, we rely on the robustness of pretrained DINOv2 features, which as shown in our experiments remain stable under the degradations present in super-resolution and deblurring tasks. An ablation in Appendix A.4 quantifies the reliability of this proxy across tasks.
@@
Extension to latent diffusion models. The proposed representation-alignment strategy applies naturally to latent diffusion models, where sampling is performed in a learned latent space rather than in pixel space. Let $z_{t}$ denote the latent state at timestep $t$. In this setting, the regularized sampling objective of eq. (11) becomes
$J_{t}(z_{t}\mid y,c_{\text{proxy}})=\log p(y\mid z_{t})+\log p_{t}(z_{t})+\lambda\,\log p(c_{\text{proxy}}\mid z_{t}).$
Following the same construction as in the pixel-space formulation, we approximate the gradient of the Repa alignment term by
$\nabla_{z_{t}}\log p(c_{\text{proxy}}\mid z_{t})\;\approx\;\nabla_{z_{t}}\sum_{n=1}^{N}\cos\!\bigl(c^{[n]}_{\text{proxy}},\,g_{\phi}(G_{2}^{[n]}(z_{t},t))\bigr),$
@@
1:Flow model $u_{\theta}=G_{1}\circ G_{2}$, measurement $y$, learning rate $\eta$, regularizer strength $\lambda$, proxy representation $c_{\text{proxy}}$
2:initialize $z_{T}\sim\mathcal{N}(0,I)$
3:for $t\in\{T,\dots,1\}$ do
@@
6:  $z_{t}\leftarrow z_{t}+\Delta t\cdot v$
7:  $z_{t-1}\leftarrow z_{t}+\eta\nabla_{z_{t}}\log p(y\mid z_{t})$
8:  $z_{t-1}\leftarrow z_{t-1}+\lambda\nabla_{z_{t}}\sum_{n=1}^{N}\cos\!\bigl(c^{[n]}_{\text{proxy}},\,g_{\phi}(G_{2}(z_{t},t)^{[n]})\bigr)$
9:end for
10:return $x_{0}$

### v2

# Align & Invert: Solving Inverse Problems with Diffusion and Flow-based Models via Representation Alignment

@@ Abstract
Enforcing alignment between the internal representations of diffusion or flow-based generative models and those of pretrained self-supervised encoders has recently been shown to provide a powerful inductive bias, improving both convergence and sample quality. In this work, we extend this idea to inverse problems, where pretrained generative models are employed as priors. We propose applying representation alignment (Repa) between diffusion or flow-based models and a DINOv2 visual encoder, to guide the reconstruction process at inference time. Although ground-truth signals are unavailable in inverse problems, we empirically show that aligning model representations of approximate target features can substantially enhance reconstruction quality and perceptual realism. We provide theoretical results showing (a) that Repa regularization can be viewed as a variational approach for minimizing a divergence measure in the DINOv2 embedding space, and (b) how under certain regularity assumptions Repa updates steer the latent diffusion states toward those of the clean image. We integrate Repa into multiple state-of-the-art inverse problem solvers, and provide extensive experiments on super-resolution, box inpainting, Gaussian deblurring, and motion deblurring confirming that our method consistently improves reconstruction quality, while also providing efficiency gains reducing the number of required discretization steps.
Machine Learning, ICML
@@ 1 Introduction
Pretrained diffusion and flow-based models, (Sohl-Dickstein et al., 2015; Ho et al., 2020; Esser et al., 2024; Lipman et al., 2023), have recently been at the heart of methods focusing on addressing various types of inverse problems. The crux of these approaches is to perform diffusion sampling, incorporating measurement information during the reverse process, with the goal to reconstruct the clean image at the final step, (Patel et al., 2024; Chung et al., 2023; Thaker et al., 2025). Undoubtedly, diffusion and flow-based models have pushed the boundaries in addressing a wealth of inverse problems, (Daras et al., 2024). However, they still struggle to produce highly detailed images particularly in cases of severe degradation and in challenging scenes, such as natural images with rich textures and complex backgrounds. These issues are further exacerbated when using latent diffusion models (Rombach et al., 2022), where the inclusion of encoder-decoder architectures introduces additional challenges during the reverse process due to nonlinearities. To mitigate these limitations, further inductive biases are needed during inference, (Rout et al., 2023; Raphaeli et al., 2025).
@@
Measurement Without REPA With REPA Ground Truth Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRepresentation AlignmentObjectiveDINOv2 EncoderRefer to captionApproximatereconstructed imageMeasurement Matching ObjectiveSiT Transformer BlockSiT Transformer BlockSiT Transformer BlockSiT Transformer BlockSiT Transformer BlockSiT Transformer BlockMLP+
@@
Figure 1: Overview of our proposed framework. Left: Box inpainting (top row) and Gaussian deblurring (bottom row) results, where adding Repa improves perceptual quality. Right: Alignment between diffusion features and pretrained DINOv2 embeddings.
@@
Recent studies have shown that diffusion models learn semantic features in their hidden states, (Li et al., 2023). The more expressive these features are the better the diffusion model performs on the generative task, (Xiang et al., 2023). Building on this insight, the seminal work of Yu et al. (2024) introduced a regularizer that aligns the internal representations of the diffusion model with those of a pretrained visual encoder, (Oquab et al., 2023). This framework, termed representation alignment (Repa), was shown to act as a strong semantic constraint leading to higher-fidelity generations and significantly faster convergence.
@@
The success of Repa has attracted considerable interest in this direction, (Wang et al., 2025b; Tian et al., 2025; Yao et al., 2025; Leng et al., 2025; Wang et al., 2025a). In particular, (Wang et al., 2025b) observed that while Repa enhances early training dynamics, its benefits diminish over time and can even degrade performance in later iterations. To address this issue, they introduced an early-stopping strategy which mitigated the late-stage degradation and further improved overall convergence speed. Other works have explored applying alignment losses directly to VAE latent spaces (Yao et al., 2025; Xu et al., 2025), finding that diffusion models trained on these regularized spaces converge faster. Despite these advances, most efforts focus on improving training efficiency, whereas relatively little attention has been paid to leveraging representation alignment at inference time, particularly for solving inverse problems. This motivates our central question:
@@
Can we apply representation alignment to benefit existing algorithms for solving inverse problems using pretrained diffusion and flow-based models?
@@
  • We introduce a general framework for solving inverse problems with diffusion and flow-based models by enforcing alignment between internal diffusion representations and DINOv2 features (Fig. 1). Despite the absence of ground truth, alignment with proxy reconstructions remains effective thanks to the robustness/invariances of DINOv2 representations.
  • We provide theoretical results that characterize the effect of the Repa regularizer: (a) its connection to a divergence measure defined in the DINOv2 feature space, and (b) under suitable regularity assumptions, its ability to steer latent diffusion states toward those associated with the clean image.
  • We demonstrate the versatility of our approach by integrating representational alignment into different existing state-of-the-art inverse problem solvers.
  • Through extensive experiments on super-resolution, Gaussian deblurring, motion deblurring, and box inpainting, we show consistent improvements over prior methods across tasks and metrics, and further demonstrate that applying Repa to existing solvers allows them to match their original performance using significantly fewer discretization steps.
@@ 2.1 Diffusion & Flow-based Models
Generative models based on denoising are a class of deep generative models that aim to learn a data distribution $p_{0}(x)$ by training on progressively corrupted versions of real samples (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021; Liu et al., 2023). These models define a) a forward (noising) process, which gradually interpolates between the data distribution and a reference distribution $p_{1}(x)$ (typically a Gaussian), and b) a reverse (denoising) process, which is trained to invert this corruption and iteratively reconstruct data from noise. Following the formulation of stochastic interpolants (Albergo et al., 2023), the time marginals of this process can be expressed as
@@
where $\alpha_{t}$ and $\sigma_{t}$ correspond to decreasing and increasing function of $t$, respectively, with boundary conditions $\alpha_{0}=\sigma_{1}=1$ and $\alpha_{1}=\sigma_{0}=0$. Throughout, we use the notation $\dot{\alpha}_{t}:=\tfrac{\mathrm{d}\alpha_{t}}{\mathrm{d}t}$ and $\dot{\sigma}_{t}:=\tfrac{\mathrm{d}\sigma_{t}}{\mathrm{d}t}$ to denote time derivatives. Lipman et al. (2023) showed that there exists a probability flow ordinary differential equation (ODE) whose solution follows the same time marginals as the process in eq. (1):
@@
where $v(x_{t},t)$ denotes the true drift field that transports $p_{0}(x)$ to $p_{1}(x)$. To enable sampling from the data distribution, one has to train a neural network $v_{\theta}(x_{t},t)$ to approximate $v(x_{t},t)$. This is done by regressing the network output toward the optimal vector field using conditional flow matching (Lipman et al., 2023; Hyvärinen and Dayan, 2005), minimizing
@@
Eq. (5) hints that the score parameterization used in diffusion models and the velocity parameterization used in flow-based models can lead to equivalent formulations of the sampling process. In practice, diffusion models correspond to a stochastic discretization of the probability flow ODE, while flow-based models correspond to its deterministic counterpart (Lai et al., 2025).
@@ 2.2 Diffusion Models with Representation Guidance
Diffusion models learn rich, discriminative features in their hidden states, which are crucial for their generative performance (Xiang et al., 2023). Nevertheless, their learned representations still fall behind on downstream tasks when compared with state-of-the-art self-supervised visual encoders, (Assran et al., 2023; Siméoni et al., 2025). This representation gap has been identified as a key bottleneck for improving the generative performance of diffusion models, as richer and more semantically structured representations have been shown to enhance sample fidelity (Yu et al., 2024). To address this, a growing line of work introduces representation guidance, which conditions diffusion models on external representations from pretrained self-supervised encoders (Li et al., 2024). Typically these approaches fall into two categories: (i) conditioning the diffusion model on external representations, either learned, (Li et al., 2024), or retrieved from a database of samples, (Blattmann et al., 2022; Hu et al., 2023; Sheynin et al., 2022); and (ii) explicitly aligning the model’s internal representations with those of large pretrained encoders, thereby encouraging more semantically meaningful feature-spaces, (Tian et al., 2025; Hu et al., 2023; Leng et al., 2025).
@@
$h_{t}=\textsc{DiffEnc}(x_{t},t),$
where $x_{t}$ denotes the noisy image and DiffEnc returns the patch-level hidden representation after the first $\ell$ blocks of the diffusion model’s transformer encoder. Repa enforces patch-wise similarity between the $D_{1}$-dimensional DINOv2 representation of the target image $x$, i.e., $f^{[n]}_{\text{DINOv2}}(x)\in\mathbb{R}^{D_{1}}$ for patches $n=1,\dots,N$, and the intermediate hidden states $h^{[n]}_{t}$, by maximizing the following objective during training:
@@
2.3 Inverse Problems with Diffusion and Flow Models
When solving inverse problems the goal is to recover the clean signal $x_{0}$ from noisy or degraded measurements $y$. In the flow- and diffusion-based framework, this is accomplished by replacing the unconditional score function in eq. (4) with the conditional score $\nabla_{x_{t}}\log p(x_{t}\mid y)$. Applying Bayes’ rule, the conditional score can be decomposed as
$\nabla_{x_{t}}\log p(x_{t}\mid y)=\nabla_{x_{t}}\log p_{t}(x_{t})+\nabla_{x_{t}}\log p(y\mid x_{t}),$
Using this decomposition, the reverse-time SDE becomes
$dx_{t}=\bigl(v(x_{t},t)-g^{2}(t)\,\nabla_{x_{t}}\log p(x_{t}\mid y)\bigr)\,dt+g(t)\,d\bar{w}_{t}.$
The unconditional score $\nabla_{x_{t}}\log p_{t}(x_{t})$ is provided by a pretrained diffusion or flow-based model, while the likelihood term $\nabla_{x_{t}}\log p(y\mid x_{t})$ enforces measurement consistency but is generally intractable, as it requires marginalizing over all possible clean signals $x_{0}$. One of the most prominent approaches is Diffusion Posterior Sampling (DPS), (Chung et al., 2023), which uses the following approximation:
$p(y\mid x_{t})\approx p\left(y\mid\hat{x}_{0}=\mathbb{E}[x_{0}\mid x_{t}]\right),$
where $\mathbb{E}[x_{0}\mid x_{t}]$ denotes the conditional expectation of $x_{0}$ given $x_{t}$. In many practical scenarios, performing the diffusion process directly in pixel space is computationally prohibitive. Latent diffusion models (LDMs) address this limitation by operating in a compressed latent space. Let $\mathcal{E}:\mathbb{R}^{k}\rightarrow\mathbb{R}^{d}$ be an encoder mapping an image $x_{0}$ to its latent representation $z_{0}=\mathcal{E}(x_{0})$, and let $\mathcal{D}$ denote the corresponding decoder such that $\hat{x}_{0}=\mathcal{D}(z_{0})\approx x_{0}$. Building on this formulation, recent work has proposed latent-space counterparts of pixel-based inverse problem solvers. For example, Diffusion Posterior Sampling (DPS) has been formulated in latent space by redefining the measurement likelihood as
$p(y\mid z_{t})\approx p\!\left(y\,\middle|\,\hat{x}_{0}=\mathcal{D}\!\big(\mathbb{E}[z_{0}\mid z_{t}]\big)\right).$
This straightforward formulation commonly referred to as Latent DPS has been observed to introduce artifacts, largely due to the nonlinearity of the decoder. To mitigate this effect, Rout et al. (2023) proposed a regularization term encouraging the latent variables to remain close to a fixed point of the encoder–decoder mapping, while Song et al. (2024) enforced data consistency during optimization followed by a projection of the latent variables back onto the noisy manifold.
@@ 3 Proposed Approach
We aim to apply Repa framework to inverse problems. However, a key challenge is that representation alignment requires access to the ground-truth signal, which is not available in this setting. To overcome this, we must choose an alternative input to the DINOv2 encoder that can stand in for the missing ground truth. In particular, we need an approximation $\bar{x}_{0}$ of the unknown $x_{0}$ that produces a proxy representation $c_{\text{proxy}}\in\mathbb{R}^{N\times D_{1}}$, where $N$ denotes the number of patch tokens and $D_{1}$ is their feature dimension. This representation should satisfy
$c_{\text{proxy}}\equiv f_{\text{DINOv2}}(\bar{x}_{0})\approx f_{\text{DINOv2}}(x_{0}).$ (7)
We then introduce a tilted distribution, (Pachebat et al., 2025), that incorporates representation alignment as an additional reward:
$\widetilde{p}(x_{t}\mid y)$ $\propto p(y\mid x_{t})\,p_{t}(x_{t})$ (8)
$\quad\exp\!\bigl(\lambda\,\textsc{Repa}(\bar{x}_{0},h_{t})\bigr),$
where $h_{t}=\textsc{DiffEnc}(x_{t},t)$ denotes the intermediate diffusion representation and $\lambda>0$ controls the strength of the alignment term. The likelihood $p(y\mid x_{t})$ enforces measurement consistency, while $p_{t}(x_{t})$ corresponds to the unconditional prior induced by the pretrained diffusion or flow model. The Repa term is defined in Eq. (6) and encourages alignment between the diffusion representation of $x_{t}$ and the proxy DINOv2 features $c_{\text{proxy}}$, thereby biasing the sampling process toward semantically consistent reconstructions.
@@
On the selection of $c_{\text{proxy}}$. To address the fact that representation alignment (Yu et al., 2024) is not directly applicable to inverse problems where the ground-truth image is unavailable at inference time, we construct $c_{\text{proxy}}$ from an approximate reconstruction. We initialize the proxy using the DINOv2 features of the available observation and, as the reverse diffusion process progresses and the reconstruction becomes more informative, we gradually replace this proxy with the features of the model’s current denoised estimate, namely $f_{\text{DINOv2}}(\mathbb{E}[x_{0}\mid x_{t}])$. In both cases, we rely on the robustness of pretrained DINOv2 features, which, as shown in our experiments, remain stable under the degradations present in super-resolution and deblurring tasks. An ablation in Appendix A.4 quantifies the reliability of this proxy across tasks.
@@
Extension to latent diffusion models. The proposed representation-alignment strategy applies naturally to latent diffusion models, where sampling is performed in a learned latent space rather than in pixel space. Let $z_{t}$ denote the latent state at timestep $t$. In this setting, the tilted distribution becomes
$\widetilde{p}(z_{t}\mid y)\propto p(y\mid z_{t})p_{t}(z_{t})\exp\left(\lambda\textsc{Repa}(\bar{x}_{0},h_{t})\right),$
With a slight abuse of notation, we use DiffEnc to denote the diffusion encoder in both pixel-space and latent-space models, and write $h_{t}=\textsc{DiffEnc}(z_{t},t)$ for latent diffusion.
@@
1:Flow or diffusion model $u_{\theta}$, measurement $y$, learning rate $\eta$, regularizer strength $\lambda$, proxy representation $c_{\text{proxy}}$, timestep $t_{\text{cutoff}}$, Decoder $\mathcal{D}$
2:Initialize $z_{T}\sim\mathcal{N}(0,I)$,
3:for $t=T,\dots,1$ do
@@
6:  $z_{t-1}\leftarrow z_{t}-\Delta t\cdot v$
7:  $z_{t-1}\leftarrow z_{t-1}+\eta\,\nabla_{z_{t}}\log p(y\mid z_{t})$
8:
9:  $z_{t-1}\leftarrow z_{t-1}+\lambda\,\nabla_{z_{t}}\sum_{n=1}^{N}\cos\!\bigl(c^{[n]}_{\text{proxy}}$,
10:     $\qquad\qquad\qquad g_{\phi}\!\big(\mathrm{DiffEnc}^{[n]}(z_{t},t))\bigr)$
11:  if $t<t_{\text{cutoff}}$ then
12:   $c_{\text{proxy}}\leftarrow f_{\mathrm{DINOv2}}\!\bigl(\mathcal{D}(\mathbb{E}[z_{0}\mid x_{t}])\bigr)$
13:  end if
14:end for
15:return $z_{0}$